# MobileFlow: A Multimodal LLM for Mobile GUI Agent

### v2

@@ 3.2.2 GUI Chain-of-Thought
4 Experiments
In this section, we begin by presenting the dataset employed for both training and testing, highlighting its enhanced compatibility with our business applications. Subsequently, we delve into the specifics of our training regimen, the environment setup, and the intricate details of our state-of-the-art algorithm implementations. We then proceed to unveil the qualitative and quantitative outcomes for the action prediction and visual question answering tasks, complemented by an ablation study that underscores the significance of our technical contributions.
4.1 Deployment Details
Dataset
@@ 4.1 Deployment Details
Training and Evaluation details
During the training of UI Encoder, we conducted fine-tuning on our 100K labeled business dataset over a span of 2 epochs. For the supervised fine-tuning phase of Qwen-vl-chat, after closely monitoring the loss function outcomes, we determined that training it for 2 epochs with a learning rate of yielded the optimal performance.
For the evaluation phase, we employed MobileAgent, making necessary adjustments to the data format and interface to ensure compatibility with our test data, all the while keeping the model parameters intact. Additionally, we directly utilized the GPT-4v interface for token prediction on our test data.
All experiments were conducted on a robust setup of 8 GPUs (Nvidia A100).
4.2 Metrics
In prior research, MobileAgent introduced a set of metrics that were effective for discrete counting with a limited number of samples—specifically, the paper referenced a case with just 10 samples. However, when scaling up to larger test datasets, these metrics fail to accurately reflect the capabilities of the proposed Large Language Model (LLM) agent. Additionally, in response to the observed issue of endpoint determination, we introduced the Endpoint Determination Rate (EDR) as a new metric to identify this problem. Endpoint determination refers to the scenario where a large language model agent does not conclude its task on the final user interface page and instead continues to generate incorrect instructions.
Furthermore, we have adopted the Whole Task Success Rate (WTSR) and Step Success Rate (SSR), metrics mentioned in previous studies, to assess the accuracy of both single-step and multi-step predictions. To elaborate, we present our metrics as follows:
Whole Task Success Rate(WTSR)
The Whole Task Success Rate (WTSR) is a metric that measures the proportion of instances where the Large Language Model (LLM) agent can successfully predict every step within a single task across the entire dataset. This metric serves as an indicator of the LLM-agent’s capability to handle tasks of varying complexities. It provides a comprehensive assessment of the agent’s performance, reflecting its ability to navigate through the full scope of a task from beginning to end successfully.
@@ 4.2 Metrics
Step Success Rate(SSR)
Given a specific intention, the Step Success Rate (SSR) measures the frequency at which the Large Language Model (LLM) agent can accurately predict each individual step within all multi-step tasks. This metric evaluates the agent’s precision in executing and forecasting each stage of a task, regardless of the overall task’s success. SSR is particularly useful for understanding the agent’s performance on individual components of complex tasks, offering insights into its ability to manage sequential decision-making processes effectively.
@@
Endpoint Determination Rate(EDR)
The Endpoint Determination Rate (EDR) is a metric that quantifies the proportion of tasks that the Large Language Model (LLM) agent successfully concludes on the final page of the entire dataset. This rate is crucial for assessing the agent’s ability to recognize when a task has reached its endpoint and to terminate its operations appropriately. A high EDR indicates that the agent is adept at identifying the conclusion of tasks and avoiding the generation of unnecessary or incorrect instructions beyond the task’s endpoint.
@@
Figure 5: Matching cases for positive and negative samples.
However, determining the correctness of a prediction is not always a clear-cut or binary decision. To address this, our proposed metrics employ a combination of methods: matching the predicted action type and calculating the Intersection over Union (IoU) of the coordinate areas to assign a true or false label to each prediction. This approach allows for a more nuanced evaluation that accounts for the complexity of the tasks and the subtleties of the predictions made by the LLM-agent. Specifically, we illustrate all the matching cases in Fig.5, providing a visual representation of how these metrics are applied to assess the accuracy of predictions.
4.3 Quantitative Results
Long Chain Tasks WTSR 0.2353 0.1765 0.1875 0.2857 - - 0.2213
SSR 0.8282 0.7665 0.8163 0.5652 - - 0.7441
EDR 0.1429 0.14 0.1574 0.1038 - - 0.1360
Middle Chain Tasks WTSR 0.7691 0.3999 0.5554 0.2308 0.2 - 0.4310
SSR 0.9634 0.9032 0.8947 0.5652 0.7547 - 0.8162
EDR 0.3241 0.423 0.1739 0.1796 0.1875 - 0.2576
Short Chain Tasks WTSR - 0.9995 - 0.4154 0.7999 0.6363 0.7128
SSR - 0.9997 - 0.4386 0.8332 0.7199 0.7479
EDR - 0.25 - 0.2015 0.2890 0.2047 0.2363
Average WTSR 0.4667 0.2917 0.3333 0.3810 0.3333 0.6363 0.4071
SSR 0.8735 0.7921 0.8341 0.5353 0.6136 0.7199 0.7280
EDR 0.3 0.2083 0.2593 0.1807 0.2450 0.2047 0.2330
For our quantitative assessment, we aimed to scrutinize the performance of MobileFlow across different business sectors. To this end, we segmented the entire testing dataset into six distinct business categories: Food Delivery Apps, Food Walk-in Apps, Insurance Apps, Medical Service Apps, Fund Selection Apps, and Gaming Apps. Furthermore, to gauge MobileFlow’s proficiency in managing tasks of varying complexities, we categorized all tasks into three levels of operational complexity: long chain tasks (those comprising more than 8 steps), middle chain tasks (those with more than 4 steps but no more than 8 steps), and short chain tasks (those with 4 steps or fewer). The comparative results are presented in Tab.1, offering a detailed view of MobileFlow’s performance metrics in comparison to existing state-of-the-art solutions.
Regarding the Endpoint Determination Rate (EDR), the initial quantitative results may appear to be on the lower side. However, it is important to understand that EDR is intrinsically linked to the Whole Task Success Rate (WTSR). A task must be both successfully executed through all its steps and correctly terminated for the endpoint determination to register it as a positive sample. Consequently, it is theoretically expected that EDR should be lower than WTSR. Our observations confirm this relationship, with tasks that WTSR identifies as successfully predicted also tending to have a higher EDR. This is particularly evident in the Medical Service App scenario, where all tasks that were successfully predicted were also properly terminated.
GPT-4v WTSR 0.1833 0.1876 0.1138 0.1428 0.2021 0.4132 0.2071
SSR 0.5716 0.4647 0.3571 0.4857 0.5012 0.6613 0.5069
Qwen-VL-Max WTSR 0.3650 0.2562 0.1643 0.2875 0.3076 0.5832 0.3273
SSR 0.7338 0.7075 0.6962 0.5112 0.2425 0.7763 0.6113
MobileFlow WTSR 0.4667 0.2917 0.3333 0.3810 0.3333 0.6363 0.4071
SSR 0.8735 0.7921 0.8341 0.5353 0.6136 0.7199 0.7280
Upon analyzing the comparative results with the state-of-the-art Large Language Model (LLM) agents, it is evident that MobileFlow demonstrates superior quantitative performance across various business sectors. Our experimental observations indicate that MobileAgent delivers optimal outcomes when faced with complex tasks. Furthermore, although the LLM integrated into MobileFlow is considerably smaller in size compared to Qwen-vl-max, the data presented in Tab.2 reveal that these two methodologies can yield competitive results relative to one another. This suggests that MobileFlow, even with a smaller LLM, can achieve enhanced outcomes post-supervised fine-tuning. The competitive results underscore the effectiveness of MobileFlow’s approach and its potential to outperform or at least rival larger models, highlighting the efficiency and robustness of our proposed method.
For the visual-question-answering tasks, the results are presented in the subsequent Tab.3. This table provides a detailed analysis of how MobileFlow and other state-of-the-art LLM-agents perform on tasks that require understanding both visual and textual information to generate accurate responses. The data in Table 3 will offer insights into the effectiveness of each method in handling the complexities of visual-question-answering, including the precision of answers and the ability to correctly interpret and respond to the questions posed alongside visual stimuli.
Table 3: Comparison with SOTA LLM-agent on VQA task
Method Recall Accuracy F-score
GPT-4v 0.6835 0.6228 0.6521
Qwen-VL-Max 0.7478 0.7064 0.7268
MobileFlow 0.7478 0.7253 0.7363
Our MobileFlow has demonstrated robust visual-question-answering (VQA) capabilities, particularly after undergoing supervised fine-tuning with our specific business data. Remarkably, it holds its ground even when compared to larger language models like Qwen-vl-max. This achievement underscores the effectiveness of our fine-tuning process and the adaptability of MobileFlow to perform well in VQA tasks, despite having a smaller model size.
4.4 Ablation Study
The Effectiveness of MoE structure
Based on the data from Tab.4, the newly proposed architecture of ViT (Vision Transformer) with MoE (Mixture of Experts) in MobileFlow has achieved significant improvements over the conventional ViT plus dense Large Language Model (LLM) structure. Specifically, there is an 4.49% enhancement in Whole Task Success Rate (WTSR) and a notable 8.17% increase in Step Success Rate (SSR). These results underscore the effectiveness of the MoE components in enhancing the performance of MobileFlow, highlighting the benefits of this innovative model structure for complex task execution and prediction accuracy.
Table 4: Effectiveness of the MoE structure
Model WTSR SSR
ViT(448px) + Dense 0.2926 0.6016
ViT(448px) + MoE 0.3375 0.6833
The Effectiveness of the UI encoder
We compared the performance of MobileFlow before and after the inclusion of the UI Encoder. Experimental results indicate that incorporating the UI Encoder as a visual branch into the MobileFlow architecture resulted in a 6.96% improvement in the WTSR metric and a 4.47% improvement in the SSR metric. This further underscores the importance of supplementing UI visual information for the final decision-making in MobileFlow.
Table 5: Effectiveness of the UI encoder
Model WTSR SSR
ViT(448px) + MoE 0.3375 0.6833
ViT(448px) + UI Encoder + MoE 0.4071 0.7280
5 Applications
5.1 Software Testing
@@
5.2 Advertisement Preview And Audit
@@
5.3 E-Commerce Operation And Monitoring
@@
6 Conclusion
In this paper, we present MobileFlow, a GUI Agent that leverages a multimodal large model and incorporates a suite of optimization techniques. MobileFlow has been designed to navigate a diverse array of intricate scenarios and business domains with proficiency. It has demonstrated its capabilities in practical applications across various sectors, including dining, healthcare, logistics, and beyond, showcasing its versatility and effectiveness.
As with many pioneering agents in the industry, MobileFlow marks an important initial step in the evolution of GUI Agents. However, MobileFlow, like its contemporaries, still faces certain limitations. It is susceptible to hallucinations, which can lead to incorrect interpretations of the GUI, and currently, it struggles with handling multiple images simultaneously. These challenges are areas for future optimization. Besides, MobileFlow’s current applications are predominantly focused on mobile apps, the future holds the promise of expansion. There is significant potential for MobileFlow to extend its reach to a variety of other devices, such as computers. In this broader context, MobileFlow could evolve into an exceptionally reliable and user-friendly intelligent assistant. It could assist humans in addressing a wide spectrum of everyday challenges, enhancing user experience, and providing valuable support across different platforms and environments.
The journey of MobileFlow, from its current state to its future potential, underscores the dynamic and rapidly evolving nature of AI and its applications. As research progresses and technology advances, MobileFlow and other similar agents are poised to become increasingly integral to our daily lives, offering innovative solutions and transforming the way we interact with digital interfaces.
Appendix A Action Space
Table 6: Action Space
Action Parameters Explanation
Click Position Click at a specified location
Long Press Position Long press at a specified location
Input Text Input the text at the current cursor position
Scroll Position List Slide along a list of positions
Drag Position List Long press, then slide along a list of positions
Wait Time Wait without performing any actions
Appendix B More Samples
@@
Figure 7: User’s instruction: Check my claim records.

### v3

@@ B.1 Prompt structure
Appendix C Evaluation Details
C.1 Positive Sample Determination
Figure 4: Matching cases for positive and negative samples.
Determining the correctness of a prediction is not always a clear-cut or binary decision. To address this, our proposed metrics employ a combination of methods: matching the predicted action type and calculating the Intersection over Union (IoU) of the coordinate areas to assign a true or false label to each prediction. This approach allows for a more nuanced evaluation that accounts for the complexity of the tasks and the subtleties of the predictions made by the LLM-agent. Specifically, we illustrate all the matching cases in Fig.4, providing a visual representation of how these metrics are applied to assess the accuracy of predictions.
C.2 Deployment Details
C.3 Dataset
@@
C.4 Training and Evaluation details
During the training of UI Encoder, we conducted fine-tuning on our 100K labeled business dataset over a span of 2 epochs. For the supervised fine-tuning phase of Qwen-vl-chat, after closely monitoring the loss function outcomes, we determined that training it for 2 epochs with a learning rate of yielded the optimal performance. For the evaluation phase, we employed MobileAgent, making necessary adjustments to the data format and interface to ensure compatibility with our test data, all the while keeping the model parameters intact. Additionally, we directly utilized the GPT-4v interface for token prediction on our test data. All experiments were conducted on a robust setup of 8 GPUs (Nvidia A100).
C.5 Definition of Metrics
Whole Task Success Rate(WTSR) The Whole Task Success Rate (WTSR) is a metric that measures the proportion of instances where the Large Language Model (LLM) agent can successfully predict every step within a single task across the entire dataset.
@@ C.5 Definition of Metrics
Step Success Rate(SSR) Given a specific intention, the Step Success Rate (SSR) measures the frequency at which the Large Language Model (LLM) agent can accurately predict each individual step within all multi-step tasks.
@@
Endpoint Determination Rate(EDR) The Endpoint Determination Rate (EDR) is a metric that quantifies the proportion of tasks that the Large Language Model (LLM) agent successfully concludes on the final page of the entire dataset.
@@
Appendix D Applications
D.1 Software Testing
@@
D.2 Advertisement Preview And Audit
@@
D.3 E-Commerce Operation And Monitoring
@@
Appendix E More Samples
Figure 5: User’s instruction: Check my claim records.